# Quadratic Chabauty Experiments on Genus 2 Bielliptic Modular Curves in the LMFDB

### v1

@@ 8. Appendix: Tables
Label Model $(D,p)$ Number Field Points
11.66.2.a.1 $y^{2}=11x^{6}+11x^{4}-7x^{2}+1$ [7,23] $\mathbb{Q}(\sqrt{-7})$ (±$(\pm\frac{\sqrt{-7}}{3},\frac{76}{27})$
16.24.2.f.1 $y^{2}=4x^{6}+6x^{4}+4x^{2}+1$ [6,5] $\mathbb{Q}(i)$ (±$(\pm i,\pm i)$
16.24.2.f.1 $y^{2}=4x^{6}+6x^{4}+4x^{2}+1$ [7,37] $\mathbb{Q}(\sqrt{-7})$ (±$(\pm\frac{2\sqrt{-7}}{7},\pm\frac{5\sqrt{-7}}{49})$
16.48.2.bm.1 $y^{2}=8x^{6}+16x^{4}+9x^{2}+1$ [5,41] $\mathbb{Q}(\sqrt{-5})$ (±$(\pm\frac{3\sqrt{-5}}{7},\pm\frac{62}{343})$
16.48.2.bx.1 $y^{2}=-x^{6}+7x^{4}-7x^{2}+1$ [1,5] $\mathbb{Q}(i)$ (±$(\pm i,\pm 4)$
16.48.2.bx.1 $y^{2}=-x^{6}+7x^{4}-7x^{2}+1$ [17,7] $\mathbb{Q}(\sqrt{-17})$ (±$(\pm\sqrt{-17},\pm 84)$
16.48.2.c.1 $y^{2}=-x^{6}-5x^{4}+5x^{2}+1$ [7,11] $\mathbb{Q}(\sqrt{-7})$ (±$(\pm\sqrt{-7},\pm 8)$
16.48.2.c.1 $y^{2}=-x^{6}-5x^{4}+5x^{2}+1$ [1,17] $\mathbb{Q}(i,\sqrt{2})$ (±$(\pm i,\pm 2\sqrt{-2})$
16.48.2.c.2 $y^{2}=-2x^{6}-10x^{4}+10x^{2}+2$ [19,11] $\mathbb{Q}(\sqrt{-19})$ (±$(\pm\frac{9\sqrt{-19}}{19},\pm\frac{680\sqrt{-19}}{361})$
16.48.2.c.2 $y^{2}=-2x^{6}-10x^{4}+10x^{2}+2$ [1,17] $\mathbb{Q}(i)$ (±$(\pm i,\pm 4i)$
16.48.2.y.1 $y^{2}=-8x^{6}+18x^{4}-8x^{2}+1$ [15,17] $\mathbb{Q}(\sqrt{2})$ (±$(\pm\frac{\sqrt{2}}{2},\pm\frac{\sqrt{2}}{2})$
18.54.2.e.1 $y^{2}=-3x^{6}+18x^{4}-3x^{2}+4$ [3,7] $\mathbb{Q}(\sqrt{-3})$ (±$(\pm\frac{\sqrt{-3}}{3},\pm\frac{8}{3}),$ (±$(\pm\sqrt{-3},\pm 16)$
20.30.2.l.1 $y^{2}=256x^{6}+113x^{4}+18x^{2}+1$ [7,11] $\mathbb{Q}(\sqrt{-7})$ (±$(\pm\frac{\sqrt{-7}}{7},\pm\frac{2\sqrt{-7}}{49})$
24.36.2.a.1 $y^{2}=8x^{6}+1$ [7,37] $\mathbb{Q}(\sqrt{-7})$ (±$(\pm\frac{2\sqrt{-7}}{7},\pm\frac{13\sqrt{-7}}{49})$
24.36.2.bo.1 $y^{2}=72x^{6}+48x^{4}+12x^{2}+1$ [10,13] $\mathbb{Q}(\sqrt{-3})$ (±$(\pm\frac{\sqrt{-3}}{3},\pm\frac{\sqrt{-3}}{3})$
24.36.2.bo.1 $y^{2}=72x^{6}+48x^{4}+12x^{2}+1$ [2,19] $\mathbb{Q}(\sqrt{-2})$ (±$(\pm\frac{\sqrt{-2}}{2},\pm\sqrt{-2})$
24.36.2.bv.1 $y^{2}=-2x^{6}+2$ [1,13] $\mathbb{Q}(i)$ (±$(\pm i,\pm 2)$
24.36.2.ev.1 $y^{2}=x^{6}-13x^{4}+39x^{2}-27$ [3,13] $\mathbb{Q}(\sqrt{-3})$ $(0,\pm 3\sqrt{-3})$
24.36.2.ev.1 $y^{2}=x^{6}-13x^{4}+39x^{2}-27$ [3,19] $\mathbb{Q}(\sqrt{-3},\sqrt{-2})$ (±$(\pm\sqrt{-3},\pm 12\sqrt{-2})$
24.36.2.gc.1 $y^{2}=-6x^{6}+6$ [1,13] $\mathbb{Q}(\sqrt{3},i)$ (±$(\pm i,\pm 2\sqrt{3})$
24.36.2.gh.1 $y^{2}=-6x^{6}-78x^{4}+78x^{2}+6$ [1,5] $\mathbb{Q}(i)$ (±$(\pm i,\pm 12i)$
24.36.2.gj.1 $y^{2}=-9x^{6}-21x^{4}+21x^{2}+9$ [1,5] $\mathbb{Q}(i,\sqrt{-6})$ (±$(\pm i,\pm 2\sqrt{-6})$
24.36.2.gk.1 $y^{2}=x^{6}+13x^{4}-13x^{2}-1$ [1,5] $\mathbb{Q}(i,\sqrt{6})$ (±$(\pm i,\pm 2\sqrt{6})$
24.36.2.gk.1 $y^{2}=x^{6}+13x^{4}-13x^{2}-1$ [1,13] $\mathbb{Q}(i)$ $(0,\pm i)$
24.36.2.gm.1 $y^{2}=6x^{6}+14x^{4}-14x^{2}-6$ [1,5] $\mathbb{Q}(i)$ (±$(\pm i,\pm 4)$
24.36.2.gm.1 $y^{2}=6x^{6}+14x^{4}-14x^{2}-6$ [11,5] $\mathbb{Q}(\sqrt{-11})$ (±$(\pm\frac{\sqrt{-11}}{5},\pm\frac{192}{125})$
24.36.2.x.1 $y^{2}=2x^{6}-2$ [1,13] $\mathbb{Q}(i)$ (±$(\pm i,\pm 2i)$
24.36.2.x.1 $y^{2}=2x^{6}-2$ [17,7] $\mathbb{Q}(\sqrt{-17})$ (±$(\pm\frac{15\sqrt{-17}}{17},\pm\frac{4774\sqrt{-17}}{289})$
24.72.2.ba.1 $y^{2}=24x^{6}+12x^{4}+18x^{2}+9$ [8,43] $\mathbb{Q}(\sqrt{-2})$ (±$(\pm\frac{\sqrt{-2}}{4},\pm\frac{21}{8})$
24.72.2.cm.1 $y^{2}=-9x^{6}+3x^{4}-3x^{2}+1$ [1,13] $\mathbb{Q}(i)$ (±$(\pm i,\pm 4)$
24.72.2.da.1 $y^{2}=-72x^{6}+12x^{4}-6x^{2}+1$ [8,43] $\mathbb{Q}(\sqrt{-2})$ (±$(\pm\frac{\sqrt{-2}}{2},\pm 4)$
24.72.2.db.1 $y^{2}=9x^{6}-6x^{4}+12x^{2}-8$ [10,13] $\mathbb{Q}(\sqrt{-10})$ (±$(\pm\frac{\sqrt{-10}}{15},\pm\frac{208\sqrt{-10}}{225})$
24.72.2.db.1 $y^{2}=9x^{6}-6x^{4}+12x^{2}-8$ [2,19] $\mathbb{Q}(\sqrt{-2})$ $(0,\pm 2\sqrt{-2}),$ (±$(\pm\sqrt{-2},\pm 8\sqrt{-2})$
24.72.2.fy.1 $y^{2}=216x^{6}+72x^{4}+12x^{2}+1$ [17,13] $\mathbb{Q}(\sqrt{-3})$ (±$(\pm\frac{\sqrt{-3}}{3},\pm\sqrt{-3})$
24.72.2.hy.1 $y^{2}=-6x^{6}+12x^{4}-12x^{2}+6$ [1,5] $\mathbb{Q}(i)$ (±$(\pm i,\pm 6)$
24.72.2.ja.1 $y^{2}=24x^{6}-24x^{4}+12x^{2}-3$ [2,11] $\mathbb{Q}(\sqrt{-2})$ (±$(\pm\frac{\sqrt{-2}}{2},\pm 3\sqrt{-2})$ (±$(\pm\frac{1}{2},\pm\frac{3\sqrt{-2}}{4})$
24.72.2.t.1 $y^{2}=-24x^{6}+12x^{4}-18x^{2}+9$ [6,31] $\mathbb{Q}(\sqrt{-6})$ (±$(\pm\frac{\sqrt{-6}}{2},\pm 12)$
24.72.2.u.1 $y^{2}=9x^{6}+9x^{4}+3x^{2}+3$ [1,13] $\mathbb{Q}(i)$ (±$(\pm 2i,\pm 21i)$
28.32.2.a.1 $y^{2}=2x^{6}+46x^{4}+30x^{2}-14$ [19,5] $\mathbb{Q}(\sqrt{-19})$ (±$(\pm\sqrt{-19},\pm 48)$
28.32.2.a.1 $y^{2}=2x^{6}+46x^{4}+30x^{2}-14$ [2,17] $\mathbb{Q}(\sqrt{-2})$ (±$(\pm\frac{3\sqrt{-2}}{2},\pm\frac{49}{2})$
28.48.2.g.1 $y^{2}=49x^{6}+70x^{4}+25x^{2}+4$ [17,13] $\mathbb{Q}(\sqrt{-17})$ (±$(\pm\frac{\sqrt{-17}}{7},\pm\frac{64}{49})$
28.48.2.g.1 $y^{2}=49x^{6}+70x^{4}+25x^{2}+4$ [67,19] $\mathbb{Q}(\sqrt{7})$ (±$(\pm\frac{\sqrt{7}}{7},\pm\frac{8\sqrt{7}}{7})$
30.30.2.b.1 $y^{2}=150x^{6}-690x^{4}+546x^{2}-6$ [19,17] $\mathbb{Q}(\sqrt{-15})$ (±$(\pm\frac{\sqrt{-15}}{5},\pm\frac{32\sqrt{-15}}{5})$
36.24.2.a.1 $y^{2}=x^{6}+108$ [15,19] $\mathbb{Q}(\sqrt{-3})$ (±$(\pm\sqrt{-3},\pm 9)$
36.54.2.i.1 $y^{2}=x^{6}+6x^{4}+9x^{2}+16$ [3,13] $\mathbb{Q}(\sqrt{-3})$ (±$(\pm\sqrt{-3},\pm 4)$
37.38.2.a.1 $y^{2}=-x^{6}-9x^{4}-11x^{2}+37$ [3,7] $\mathbb{Q}(\sqrt{-3})$ (±$(\pm\sqrt{-3},\pm 4)$, (±$(\pm 3\sqrt{-3},\pm 116)$
37.38.2.a.1 $y^{2}=-x^{6}-9x^{4}-11x^{2}+37$ [7,11] $\mathbb{Q}(\sqrt{-7})$ (±$(\pm\sqrt{-7},\pm 4$) (±$(\pm\frac{\sqrt{-7}}{3},\pm\frac{172}{27})$
38.40.2.b.1 $y^{2}=4x^{6}+64x^{4}+304x^{2}+361$ [3,13] $\mathbb{Q}(\sqrt{-3})$ (±$(\pm\frac{19\sqrt{-3}}{12},\pm\frac{437\sqrt{-3}}{288})$
40.30.2.e.1 $y^{2}=25x^{6}-115x^{4}+91x^{2}-1$ [1,13] $\mathbb{Q}(i)$ $(0,\pm i)$
40.30.2.f.1 $y^{2}=x^{6}-91x^{4}+115x^{2}-25$ [1,13] $\mathbb{Q}(i)$ $(0,\pm 5i)$
45.54.2.c.1 $y^{2}=5x^{6}+21x^{4}-9x^{2}-1$ [1,13] $\mathbb{Q}(i)$ $(0,\pm i)$
48.48.2.ez.1 $y^{2}=3x^{6}-21x^{4}+21x^{2}-3$ [1,13] $\mathbb{Q}(\sqrt{-3},i)$ (±$(\pm i,\pm 4\sqrt{-3})$
56.48.2.l.1 $y^{2}=2x^{6}+25x^{4}+140x^{2}+196$ [17,13] $\mathbb{Q}(\sqrt{14})$ (±$(\pm\sqrt{14},\pm 112)$
60.30.2.p.1 $y^{2}=-3x^{6}-5x^{4}-225x^{2}-3375$ [17,7] $\mathbb{Q}(\sqrt{-5})$ (±$(\pm\sqrt{-5},\pm 20\sqrt{-5})$
60.30.2.p.1 $y^{2}=-3x^{6}-5x^{4}-225x^{2}-3375$ [15,19] $\mathbb{Q}(\sqrt{-15})$ (±$(\pm\frac{\sqrt{-15}}{6},\pm\frac{355\sqrt{-15}}{24})$
60.48.2.f.1 $y^{2}=9x^{6}+60x^{4}+25x^{2}-250$ [13,17] $\mathbb{Q}(\sqrt{-15})$ (±$(\pm\frac{\sqrt{-15}}{3},\pm\frac{10\sqrt{-15}}{3})$
70.30.2.b.1 $y^{2}=x^{6}-98x^{4}+833x^{2}-10976$ [7,11] $\mathbb{Q}(\sqrt{-7})$ (±$(\pm\sqrt{-7},\pm 56\sqrt{-7})$

### v2

@@ 8. Appendix: Tables
Label Model $y^{2}=f(x)$ $(D,p)$ Number Field Points
11.66.2.a.1 $11x^{6}+11x^{4}-7x^{2}+1$ [7,23] $\mathbb{Q}(\sqrt{-7})$ $(\pm\frac{\sqrt{-7}}{3},\frac{76}{27})$
16.24.2.f.1 $4x^{6}+6x^{4}+4x^{2}+1$ [6,5] $\mathbb{Q}(i)$ $(\pm i,\pm i)$
16.24.2.f.1 $4x^{6}+6x^{4}+4x^{2}+1$ [7,37] $\mathbb{Q}(\sqrt{-7})$ $(\pm\frac{2\sqrt{-7}}{7},\pm\frac{5\sqrt{-7}}{49})$
16.48.2.bm.1 $8x^{6}+16x^{4}+9x^{2}+1$ [5,41] $\mathbb{Q}(\sqrt{-5})$ $(\pm\frac{3\sqrt{-5}}{7},\pm\frac{62}{343})$
16.48.2.bx.1 $x^{6}+7x^{4}-7x^{2}+1$ [1,5] $\mathbb{Q}(i)$ $(\pm i,\pm 4)$
16.48.2.bx.1 $x^{6}+7x^{4}-7x^{2}+1$ [17,7] $\mathbb{Q}(\sqrt{-17})$ $(\pm\sqrt{-17},\pm 84)$
16.48.2.c.1 $x^{6}-5x^{4}+5x^{2}+1$ [7,11] $\mathbb{Q}(\sqrt{-7})$ $(\pm\sqrt{-7},\pm 8)$
16.48.2.c.1 $x^{6}-5x^{4}+5x^{2}+1$ [1,17] $\mathbb{Q}(i,\sqrt{2})$ $(\pm i,\pm 2\sqrt{-2})$
16.48.2.c.2 $2x^{6}-10x^{4}+10x^{2}+2$ [19,11] $\mathbb{Q}(\sqrt{-19})$ $(\pm\frac{9\sqrt{-19}}{19},\pm\frac{680\sqrt{-19}}{361})$
16.48.2.c.2 $2x^{6}-10x^{4}+10x^{2}+2$ [1,17] $\mathbb{Q}(i)$ $(\pm i,\pm 4i)$
16.48.2.y.1 $8x^{6}+18x^{4}-8x^{2}+1$ [15,17] $\mathbb{Q}(\sqrt{2})$ $(\pm\frac{\sqrt{2}}{2},\pm\frac{\sqrt{2}}{2})$
18.54.2.e.1 $3x^{6}+18x^{4}-3x^{2}+4$ [3,7] $\mathbb{Q}(\sqrt{-3})$ $(\pm\frac{\sqrt{-3}}{3},\pm\frac{8}{3}),$ $(\pm\sqrt{-3},\pm 16)$
20.30.2.l.1 $256x^{6}+113x^{4}+18x^{2}+1$ [7,11] $\mathbb{Q}(\sqrt{-7})$ $(\pm\frac{\sqrt{-7}}{7},\pm\frac{2\sqrt{-7}}{49})$
24.36.2.a.1 $8x^{6}+1$ [7,37] $\mathbb{Q}(\sqrt{-7})$ $(\pm\frac{2\sqrt{-7}}{7},\pm\frac{13\sqrt{-7}}{49})$
24.36.2.bo.1 $72x^{6}+48x^{4}+12x^{2}+1$ [10,13] $\mathbb{Q}(\sqrt{-3})$ $(\pm\frac{\sqrt{-3}}{3},\pm\frac{\sqrt{-3}}{3})$
24.36.2.bo.1 $72x^{6}+48x^{4}+12x^{2}+1$ [2,19] $\mathbb{Q}(\sqrt{-2})$ $(\pm\frac{\sqrt{-2}}{2},\pm\sqrt{-2})$
24.36.2.bv.1 $2x^{6}+2$ [1,13] $\mathbb{Q}(i)$ $(\pm i,\pm 2)$
24.36.2.ev.1 $x^{6}-13x^{4}+39x^{2}-27$ [3,13] $\mathbb{Q}(\sqrt{-3})$ $(0,\pm 3\sqrt{-3})$
24.36.2.ev.1 $x^{6}-13x^{4}+39x^{2}-27$ [3,19] $\mathbb{Q}(\sqrt{-3},\sqrt{-2})$ $(\pm\sqrt{-3},\pm 12\sqrt{-2})$
24.36.2.gc.1 $6x^{6}+6$ [1,13] $\mathbb{Q}(\sqrt{3},i)$ $(\pm i,\pm 2\sqrt{3})$
24.36.2.gh.1 $6x^{6}-78x^{4}+78x^{2}+6$ [1,5] $\mathbb{Q}(i)$ $(\pm i,\pm 12i)$
24.36.2.gj.1 $9x^{6}-21x^{4}+21x^{2}+9$ [1,5] $\mathbb{Q}(i,\sqrt{-6})$ $(\pm i,\pm 2\sqrt{-6})$
24.36.2.gk.1 $x^{6}+13x^{4}-13x^{2}-1$ [1,5] $\mathbb{Q}(i,\sqrt{6})$ $(\pm i,\pm 2\sqrt{6})$
24.36.2.gk.1 $x^{6}+13x^{4}-13x^{2}-1$ [1,13] $\mathbb{Q}(i)$ $(0,\pm i)$
24.36.2.gm.1 $6x^{6}+14x^{4}-14x^{2}-6$ [1,5] $\mathbb{Q}(i)$ $(\pm i,\pm 4)$
24.36.2.gm.1 $6x^{6}+14x^{4}-14x^{2}-6$ [11,5] $\mathbb{Q}(\sqrt{-11})$ $(\pm\frac{\sqrt{-11}}{5},\pm\frac{192}{125})$
24.36.2.x.1 $2x^{6}-2$ [1,13] $\mathbb{Q}(i)$ $(\pm i,\pm 2i)$
24.36.2.x.1 $2x^{6}-2$ [17,7] $\mathbb{Q}(\sqrt{-17})$ $(\pm\frac{15\sqrt{-17}}{17},\pm\frac{4774\sqrt{-17}}{289})$
24.72.2.ba.1 $24x^{6}+12x^{4}+18x^{2}+9$ [8,43] $\mathbb{Q}(\sqrt{-2})$ $(\pm\frac{\sqrt{-2}}{4},\pm\frac{21}{8})$
24.72.2.cm.1 $9x^{6}+3x^{4}-3x^{2}+1$ [1,13] $\mathbb{Q}(i)$ $(\pm i,\pm 4)$
24.72.2.da.1 $72x^{6}+12x^{4}-6x^{2}+1$ [8,43] $\mathbb{Q}(\sqrt{-2})$ $(\pm\frac{\sqrt{-2}}{2},\pm 4)$
24.72.2.db.1 $9x^{6}-6x^{4}+12x^{2}-8$ [10,13] $\mathbb{Q}(\sqrt{-10})$ $(\pm\frac{\sqrt{-10}}{15},\pm\frac{208\sqrt{-10}}{225})$
24.72.2.db.1 $9x^{6}-6x^{4}+12x^{2}-8$ [2,19] $\mathbb{Q}(\sqrt{-2})$ $(0,\pm 2\sqrt{-2}),$ $(\pm\sqrt{-2},\pm 8\sqrt{-2})$
24.72.2.fy.1 $216x^{6}+72x^{4}+12x^{2}+1$ [17,13] $\mathbb{Q}(\sqrt{-3})$ $(\pm\frac{\sqrt{-3}}{3},\pm\sqrt{-3})$
24.72.2.hy.1 $6x^{6}+12x^{4}-12x^{2}+6$ [1,5] $\mathbb{Q}(i)$ $(\pm i,\pm 6)$
24.72.2.ja.1 $24x^{6}-24x^{4}+12x^{2}-3$ [2,11] $\mathbb{Q}(\sqrt{-2})$ $(\pm\frac{\sqrt{-2}}{2},\pm 3\sqrt{-2})$ $(\pm\frac{1}{2},\pm\frac{3\sqrt{-2}}{4})$
24.72.2.t.1 $24x^{6}+12x^{4}-18x^{2}+9$ [6,31] $\mathbb{Q}(\sqrt{-6})$ $(\pm\frac{\sqrt{-6}}{2},\pm 12)$
24.72.2.u.1 $9x^{6}+9x^{4}+3x^{2}+3$ [1,13] $\mathbb{Q}(i)$ $(\pm 2i,\pm 21i)$
28.32.2.a.1 $2x^{6}+46x^{4}+30x^{2}-14$ [19,5] $\mathbb{Q}(\sqrt{-19})$ $(\pm\sqrt{-19},\pm 48)$
28.32.2.a.1 $2x^{6}+46x^{4}+30x^{2}-14$ [2,17] $\mathbb{Q}(\sqrt{-2})$ $(\pm\frac{3\sqrt{-2}}{2},\pm\frac{49}{2})$
28.48.2.g.1 $49x^{6}+70x^{4}+25x^{2}+4$ [17,13] $\mathbb{Q}(\sqrt{-17})$ $(\pm\frac{\sqrt{-17}}{7},\pm\frac{64}{49})$
28.48.2.g.1 $49x^{6}+70x^{4}+25x^{2}+4$ [67,19] $\mathbb{Q}(\sqrt{7})$ $(\pm\frac{\sqrt{7}}{7},\pm\frac{8\sqrt{7}}{7})$
30.30.2.b.1 $150x^{6}-690x^{4}+546x^{2}-6$ [19,17] $\mathbb{Q}(\sqrt{-15})$ $(\pm\frac{\sqrt{-15}}{5},\pm\frac{32\sqrt{-15}}{5})$
36.24.2.a.1 $x^{6}+108$ [15,19] $\mathbb{Q}(\sqrt{-3})$ $(\pm\sqrt{-3},\pm 9)$
36.54.2.i.1 $x^{6}+6x^{4}+9x^{2}+16$ [3,13] $\mathbb{Q}(\sqrt{-3})$ $(\pm\sqrt{-3},\pm 4)$
37.38.2.a.1 $x^{6}-9x^{4}-11x^{2}+37$ [3,7] $\mathbb{Q}(\sqrt{-3})$ $(\pm\sqrt{-3},\pm 4)$, $(\pm 3\sqrt{-3},\pm 116)$
37.38.2.a.1 $x^{6}-9x^{4}-11x^{2}+37$ [7,11] $\mathbb{Q}(\sqrt{-7})$ $(\pm\sqrt{-7},\pm 4$) $(\pm\frac{\sqrt{-7}}{3},\pm\frac{172}{27})$
38.40.2.b.1 $4x^{6}+64x^{4}+304x^{2}+361$ [3,13] $\mathbb{Q}(\sqrt{-3})$ $(\pm\frac{19\sqrt{-3}}{12},\pm\frac{437\sqrt{-3}}{288})$
40.30.2.e.1 $25x^{6}-115x^{4}+91x^{2}-1$ [1,13] $\mathbb{Q}(i)$ $(0,\pm i)$
40.30.2.f.1 $x^{6}-91x^{4}+115x^{2}-25$ [1,13] $\mathbb{Q}(i)$ $(0,\pm 5i)$
45.54.2.c.1 $5x^{6}+21x^{4}-9x^{2}-1$ [1,13] $\mathbb{Q}(i)$ $(0,\pm i)$
48.48.2.ez.1 $3x^{6}-21x^{4}+21x^{2}-3$ [1,13] $\mathbb{Q}(\sqrt{-3},i)$ $(\pm i,\pm 4\sqrt{-3})$
56.48.2.l.1 $2x^{6}+25x^{4}+140x^{2}+196$ [17,13] $\mathbb{Q}(\sqrt{14})$ $(\pm\sqrt{14},\pm 112)$
60.30.2.p.1 $3x^{6}-5x^{4}-225x^{2}-3375$ [17,7] $\mathbb{Q}(\sqrt{-5})$ $(\pm\sqrt{-5},\pm 20\sqrt{-5})$
60.30.2.p.1 $3x^{6}-5x^{4}-225x^{2}-3375$ [15,19] $\mathbb{Q}(\sqrt{-15})$ $(\pm\frac{\sqrt{-15}}{6},\pm\frac{355\sqrt{-15}}{24})$
60.48.2.f.1 $9x^{6}+60x^{4}+25x^{2}-250$ [13,17] $\mathbb{Q}(\sqrt{-15})$ $(\pm\frac{\sqrt{-15}}{3},\pm\frac{10\sqrt{-15}}{3})$
70.30.2.b.1 $x^{6}-98x^{4}+833x^{2}-10976$ [7,11] $\mathbb{Q}(\sqrt{-7})$ $(\pm\sqrt{-7},\pm 56\sqrt{-7})$